# A Spatio-Temporal Machine Learning Model for Mortgage Credit Risk: Default Probabilities and Loan Portfolios

### v2

@@ 1 Introduction
Accurately modeling credit default risk is a challenging and critical task for financial institutions, investors, regulators, and policy makers. Traditionally, linear models such as linear discriminant analysis, logistic regression, and linear discrete hazard models have been widely used to model bankruptcy probabilities in corporate credit risk research (Altman, 1968; Zmijewski, 1984; Shumway, 2001). These linear models are also frequently applied in the retail credit risk literature (Crook et al., 2007; Thomas et al., 2017) for predicting default probabilities of various loan types, including mortgage loans (Agarwal et al., 2012), credit card loans (Bellotti and Crook, 2009b, a, 2013), and personal loans (Banasik et al., 1999; Malik and Thomas, 2010).
@@
For modeling corporate credit risk, linear hazard models have been extended to account for temporal frailty correlation that cannot be captured by observable predictor variables (Duffie et al., 2009; Koopman et al., 2011). Similarly, for loans associated with a specific spatial location, such as mortgages and loans to small and medium-sized enterprises (SMEs), spatial dependence likely exists among loans that cannot be fully explained by observable predictor variables. Specifically, for mortgage credit risk, there are the following potential reasons for the presence of frailty correlation. First, local weather extremes such as wildfires (Issler et al., 2024) and hurricanes (Kousky et al., 2020; Calabrese et al., 2024) have an impact on mortgage credit risk that leads to correlations among defaults. Furthermore, local economic conditions that often cannot be measured using observable predictor variables likely impact default risks. For instance, regions heavily reliant on specific industries are vulnerable to concentrated job losses and income reductions during industry downturns. This can lead to widespread local defaults as borrowers face diminished ability to make payments. In addition, house prices are often locally interconnected, and when local house prices fall, many borrowers in an area can enter negative equity. This can lead to foreclosure contagion which further lowers property values. Harding et al. (2009) show that neighborhood foreclosures have the greatest contagion effect on house prices within a 300-foot radius. This highlights the importance of accounting for local dependencies between properties in immediate neighborhoods.
@@
To account for spatial frailty correlation, several linear models have been proposed in the SME credit risk literature (Fernandes and Artes, 2016; Agosto et al., 2019; Calabrese et al., 2019; Medina-Olivares et al., 2022). Recently, Berloco et al. (2023) introduced a model accounting for spatio-temporal correlation among firms while retaining a linear functional form for the predictor variables. Early approaches to account for spatial dependence in the mortgage credit risk literature include the use of independent and identically distributed frailty variables for metropolitan statistical areas (MSAs) (Kau et al., 2011) and ZIP code concentration variables (Agarwal et al., 2012). Zhu and Pace (2014) demonstrate that spatial probit models offer superior prediction accuracy for mortgage defaults compared to independent probit models. Calabrese and Crook (2020) extended this spatial methodology to linear hazard models. Further research on spatial frailty correlation for mortgage credit risk include Babii et al. (2019); Calabrese et al. (2024), and Medina-Olivares et al. (2023b) who introduce a linear spatio-temporal model to predict mortgage prepayment. Additional prior works on spatial frailty modeling outside the credit risk literature include Li and Ryan (2002), Banerjee and Carlin (2003), Hanson et al. (2012), and Čížek et al. (2017).
@@
Prior research has also investigated the effectiveness of machine learning methods like random forests, tree-boosting, and deep neural networks in modeling corporate credit risk (Barboza et al., 2017; Zieba et al., 2016; Sigrist and Hirnschall, 2019; Sigrist and Leuenberger, 2023; Cheraghali and Molnár, 2024) and retail credit risk (Lessmann et al., 2015; Fitzpatrick and Mues, 2016; Xia et al., 2017; Thomas et al., 2017). These non-linear machine learning methods often outperform traditional linear models in prediction accuracy by capturing complex interactions and non-linear effects in the predictor variables. For instance, Fitzpatrick and Mues (2016) find that tree-boosting is preferred over logistic regression for predicting mortgage defaults. To the best of our knowledge, there exists no prior work that uses state-of-the-art machine learning models and explicitly accounts for spatio-temporal frailty correlation for modeling credit default risk in general and mortgage risk in particular.
@@ 2.1 Notation and default probabilities
$P\left(\tau_{i}=t_{k+1i}\middle|\tau_{i}>t_{ki}\right)=\left(1+e^{-X_{ki}^{T}\beta}\right)^{-1},\leavevmode\nobreak\ \leavevmode\nobreak\ \beta\in\mathbb{R}^{p}.$ (1)
@@ 2.2 Accounting for spatial and spatio-temporal frailty correlation
where $f(\cdot)$ is a link function such as $f(x)=\left(1+e^{-x}\right)^{-1}$, $F(\cdot)$ is a function $F:\mathbb{R}^{p}\mapsto\mathbb{R}$, and the latent variable $b(\cdot,\cdot)$ is a zero-mean Gaussian process (Rasmussen and Williams, 2006) that accounts for spatial or spatio-temporal frailty correlation. Both $b(\cdot,\cdot)$ and $F(\cdot)$ are specified in the following.
For the latent frailty process $b(t,s)$, we consider two cases: a spatial model and a spatio-temporal model. In the spatial model, the Gaussian process $b(t,s)$ varies over space only and is constant over time. It is defined by a spatial covariance function $\text{Cov}(b(t,s),b(t^{\prime},s^{\prime}))=c_{\theta}(s,s^{\prime})$, $s,s^{\prime}\in\mathcal{D}$, which depends on a set of parameters $\theta\in\Theta\subset\mathbb{R}^{q}$. For more information on covariance functions in general, see Rasmussen and Williams (2006, Chapter 4), and for spatial covariance functions in particular, see Cressie (2015, Section 2.3.1). In the application below, we use a Matérn covariance function:
$c_{\theta}(s,s^{\prime})=\sigma_{1}^{2}\frac{2^{\nu-1}}{\Gamma(\nu)}\left(\frac{\sqrt{2\nu}||s-s^{\prime}||}{\rho_{s}}\right)^{\nu}K_{\nu}\left(\frac{\sqrt{2\nu}||s-s^{\prime}||}{\rho_{s}}\right),$ (4)
where $\sigma_{1}^{2}$ is a marginal variance parameter, $\rho_{s}$ is a spatial range, or scale, parameter, $\nu$ is a smoothness parameter, $\Gamma(\cdot)$ is the Gamma function, and $K_{\nu}(\cdot)$ is the modified Bessel function of the second kind. For spatio-temporal models, the Gaussian process varies over both space and time, and it is defined by a space-time covariance function $\text{Cov}(b(t,s),b(t^{\prime},s^{\prime}))=c_{\theta}((t,s),(t^{\prime},s^{\prime}))$, $(t,s),(t^{\prime},s^{\prime})\in[0,T]\times\mathcal{D}\subset\mathbb{R}^{3}$. For more information on space-time covariance functions, we refer to Cressie and Wikle (2011, Section 6.1). In our application, we use an anisotropic spatio-temporal Matérn covariance function:
$c_{\theta}((t,s),(t^{\prime},s^{\prime}))=\sigma_{1}^{2}\frac{2^{\nu-1}}{\Gamma(\nu)}\left(\sqrt{2\nu}||A^{-1}((t,s)-(t^{\prime},s^{\prime}))||\right)^{\nu}K_{\nu}\left(\sqrt{2\nu}||A^{-1}((t,s)-(t^{\prime},s^{\prime}))||\right),$ (5)
where $A=\text{diag}(\rho_{t},\rho_{s},\rho_{s})$ is a diagonal matrix containing a temporal and spatial range parameters $\rho_{t}$ and $\rho_{s}$, respectively. We mainly choose this space-time covariance function since it is a parsimonious model that allows for computationally efficient inference when using a Vecchia approximation and a correlation-based neighbor selection approach (Kang and Katzfuss, 2023), see below in Section 3.3 for more details.
@@
[…]
@@ 2.3 A tree-boosted spatio-temporal Gaussian process model
In the following, we show how to relax the linearity assumption for the fixed effects predictor variable function $F(\cdot)$ using tree-boosting in models with latent Gaussian processes. Tree-boosting (Friedman et al., 2000; Friedman, 2001; Bühlmann and Hothorn, 2007; Sigrist, 2021) is a machine learning technique that often achieves superior prediction accuracy on tabular data sets (Nielsen, 2016; Shwartz-Ziv and Armon, 2022; Januschowski et al., 2022; Grinsztajn et al., 2022). For instance, recently Grinsztajn et al. (2022) showed that tree-boosting outperforms random forest and various state-of-the-art deep neural networks on a large collection of data sets. Sigrist and Leuenberger (2023) and Cheraghali and Molnár (2024) find similar results for credit risk data.
We assume the model specified in (3) in the subsequent paragraph. In addition, we assume that $F(\cdot)$ is a function in a normed function space $\mathcal{H}$ that is the linear span of a set $\mathcal{S}$ of base learners $f_{j}(\cdot):\mathbb{R}^{p}\rightarrow\mathbb{R}$, which consist of regression trees (Breiman et al., 1984) in this article. Our goal is then to find a joint minimizer for $F(\cdot)\in\mathcal{H}$ and $\theta\in\Theta$ of the functional obtained when plugging $F(\cdot)$ and $\theta$ into the negative log-likelihood:
@@
where $F(X)$ is the row-wise evaluation of $F(\cdot)$ at $X\in\mathbb{R}^{n\times p}$, which is the matrix containing predictor variables for all observations, $n=\sum_{i=1}^{N}n_{i}$, and $\mathcal{L}(F,\theta)$ is given in (6). The minimization of the empirical risk functional in (7) is done iteratively using the latent Gaussian model boosting (LaGaBoost) algorithm given in Algorithm 1, which performs a form of functional gradient descent. In detail, this algorithm iterates between, first, finding a maximum for $\theta$ of $\mathcal{L}(F_{m-1},\theta)$ conditional on the current estimate $F_{m-1}(\cdot)$ and, second, updating the ensemble of trees $F(\cdot)$ using one functional gradient descent step given the current estimates $\theta_{m}$ and $F_{m-1}(\cdot)$. Specifically, the boosting update $f_{m}(\cdot)$ in iteration $m$ is given by the least squares approximation to the vector obtained when evaluating the negative functional gradient of the functional defined in (7) at $(F_{m-1}(\cdot),I_{X_{ki}}(\cdot))$, where $I_{X_{ki}}(\cdot)$ are indicator functions which equal $1$ at $X_{ki}$ and 00 otherwise. Equivalently, $f_{m}(\cdot)$ is the minimizer of a first-order functional Taylor approximation of the functional in (7) with $F(\cdot)=F_{m-1}(\cdot)+f(\cdot)$ around $F_{m-1}(\cdot)$ with an $L^{2}$ penalty on $f(\cdot)$ evaluated at $(X_{ki})$. For more details on the LaGaBoost algorithm, e.g., the calculation of functional gradients; see Sigrist (2022, 2023).
@@
In order that we can do estimation and prediction in a computationally feasible manner on large data sets in practice, we need to apply some approximations for both the linear Gaussian process and the tree-boosted Gaussian process models. In the following, we describe these approximations. First, the integral in (6) cannot be calculated in closed form, and we approximate it using a Laplace approximation. Laplace approximations are computationally efficient, have asymptotic convergence guarantees, and are accurate for large data sets; see, e.g., Kündig and Sigrist (2025). Furthermore, to ensure that computations with Gaussian processes scale to large data sets, we use Vecchia approximations (Vecchia, 1988; Datta et al., 2016; Katzfuss and Guinness, 2021) for the latent Gaussian process $b$. In spatial statistics, Vecchia approximations have recently “emerged as a leader among the sea of approximations” (Guinness, 2021) and are often considered as “the most promising class of approximations” (Kang and Katzfuss, 2023; Rambelli and Sigrist, 2025). In brief, Vecchia approximations calculate an approximate sparse reverse Cholesky factor of the precision matrix of the latent Gaussian process $b$ by using an ordered conditional approximation for the density of the Gaussian process. In doing so, every row of this sparse Cholesky factor contains maximally $\tilde{m}$ non-zero entries. For prediction, a Vecchia approximation is applied to the joint distribution of a latent Gaussian process at the training and prediction points. Posterior predictive distributions for the latent Gaussian process are then obtained as conditional distributions of this approximated joint distribution. Further, predictive probabilities for the observable response variables are calculated by numerically integrating over the posterior predictive distribution of the latent Gaussian variable using, e.g., adaptive Gauss-Hermite quadrature (Liu and Pierce, 1994). We describe the construction of Vecchia approximations for the latent Gaussian process in more detail in Appendix A.1. For more information on Vecchia-Laplace approximations, including estimation and prediction, see Sigrist (2023) and Kündig and Sigrist (2025). Additionally, for increased computational efficiency, we use the iterative methods of Kündig and Sigrist (2025) for estimation and prediction with Vecchia-Laplace approximations instead of Cholesky decompositions.
@@ 3.2 Predictor variables
From the data provided by Freddie Mac, we construct several static and temporally varying predictor variables. The latter predictor variables are based on information available at the beginning of each year and include the age of the loan in months, the current loan-to-value ratio, and the interest rate spread. The current loan-to-value ratio is calculated by dividing the current unpaid principal balance by the appraised value of the property at the time of origination. To calculate the interest rate spread, we follow Calabrese et al. (2024) and subtract from the loan interest rates the average interest rate currently issued by Freddie Mac for 30-year fixed-rate mortgages. This average interest rate is retrieved from the Federal Reserve Bank of St. Louishttps://fred.stlouisfed.org/series/MORTGAGE30US. Static predictor variables include the borrower’s credit score at the time when the mortgage originated, additional financial indicators, and several characteristics about the property and the mortgage. These predictor variables are commonly chosen when Freddie Mac’s loan-level data set is used in the credit risk literature; see, e.g., Hu and Zhou (2019) and Medina-Olivares et al. (2023a). In addition, we include a number of macroeconomic predictor variables at state and national levels, which vary over time and are primarily sourced from the U.S. Bureau of Economic Analysishttps://bea.gov. A description of all predictor variables is provided in Table 1. Missing values for categorical predictor variables are imputed using the most frequent category, and numeric predictor variables are imputed using the mean. Summary statistics of the numeric and categorical variables are provided in Tables 4 and 5, respectively, in Appendix A.3. For all linear models, we additionally include an intercept term in the predictor variables $X_{ki}$.

### v3

@@ 1 Introduction
Accurately modeling credit default risk is a challenging and critical task for financial institutions, investors, regulators, and policy makers. Traditionally, linear models such as linear discriminant analysis, logistic regression, and linear discrete hazard models have been widely used to model bankruptcy probabilities in corporate credit risk research (altman1968financial; zmijewski1984methodological; shumway2001forecasting). These linear models are also frequently applied in the retail credit risk literature (crook2007recent; thomas2017credit) for predicting default probabilities of various loan types, including mortgage loans (agarwal2012thy), credit card loans (bellotti2009support; bellotti2009credit; bellotti2013forecasting), and personal loans (banasik1999not; malik2010modelling).
@@
For modeling corporate credit risk, linear hazard models have been extended to account for temporal frailty correlation that cannot be captured by observable predictor variables (duffie2009frailty; koopman2011modeling). Similarly, for loans associated with a specific spatial location, such as mortgages and loans to small and medium-sized enterprises (SMEs), spatial dependence likely exists among loans that cannot be fully explained by observable predictor variables. Specifically, for mortgage credit risk, there are the following potential reasons for the presence of frailty correlation. First, local weather extremes such as wildfires (issler2024housing) and hurricanes (kousky2020flood; calabrese2024impacts) have an impact on mortgage credit risk that leads to correlations among defaults. Furthermore, local economic conditions that often cannot be measured using observable predictor variables likely impact default risks. For instance, regions heavily reliant on specific industries are vulnerable to concentrated job losses and income reductions during industry downturns. This can lead to widespread local defaults as borrowers face diminished ability to make payments. In addition, house prices are often locally interconnected, and when local house prices fall, many borrowers in an area can enter negative equity. This can lead to foreclosure contagion which further lowers property values. harding2009contagion show that neighborhood foreclosures have the greatest contagion effect on house prices within a 300-foot radius. This highlights the importance of accounting for local dependencies between properties in immediate neighborhoods.
@@
To account for spatial frailty correlation, several linear models have been proposed in the SME credit risk literature (fernandes2016spatial; agosto2019spatial; calabrese2019birds; medina2022spatial). Recently, berloco2023forecasting introduced a model accounting for spatio-temporal correlation among firms while retaining a linear functional form for the predictor variables. Early approaches to account for spatial dependence in the mortgage credit risk literature include the use of independent and identically distributed frailty variables for metropolitan statistical areas (MSAs) (kau2011analysis) and ZIP code concentration variables (agarwal2012thy). zhu2014modeling demonstrate that spatial probit models offer superior prediction accuracy for mortgage defaults compared to independent probit models. calabrese2020spatial extended this spatial methodology to linear hazard models. Further research on spatial frailty correlation for mortgage credit risk include babii2019commercial; calabrese2024impacts, and medina2023spacetime who introduce a linear spatio-temporal model to predict mortgage prepayment. Additional prior works on spatial frailty modeling outside the credit risk literature include li2002modeling, banerjee2003semiparametric, hanson2012bayesian, and vcivzek2017neighbours.
@@
Prior research has also investigated the effectiveness of machine learning methods like random forests, tree-boosting, and deep neural networks in modeling corporate credit risk (barboza2017machine; zikeba2016; sigrist2019grabit; sigrist2023machine; cheraghali2025sme) and retail credit risk (lessmann2015benchmarking; fitzpatrick2016empirical; xia2017boosted; thomas2017credit). These non-linear machine learning methods often outperform traditional linear models in prediction accuracy by capturing complex interactions and non-linear effects in the predictor variables. For instance, fitzpatrick2016empirical find that tree-boosting is preferred over logistic regression for predicting mortgage defaults. To the best of our knowledge, there exists no prior work that uses state-of-the-art machine learning models and explicitly accounts for spatio-temporal frailty correlation for modeling credit default risk in general and mortgage risk in particular.
@@ 2.1 Notation and default probabilities
$P\left(\tau_{i}=t_{k+1i}\middle|\tau_{i}>t_{ki}\right)=\left(1+e^{-X_{ki}^{T}\beta}\right)^{-1},\penalty 10000\ \penalty 10000\ \beta\in\mathbb{R}^{p}.$ (1)
@@ 2.2 Accounting for spatial and spatio-temporal frailty correlation
where $f(\cdot)$ is a link function such as $f(x)=\left(1+e^{-x}\right)^{-1}$, $F(\cdot)$ is a function $F:\mathbb{R}^{p}\mapsto\mathbb{R}$, and the latent variable $b(\cdot,\cdot)$ is a zero-mean Gaussian process (williams2006gaussian) that accounts for spatial or spatio-temporal frailty correlation. Both $b(\cdot,\cdot)$ and $F(\cdot)$ are specified in the following.
For the latent frailty process $b(t,s)$, we consider two cases: a spatial model and a spatio-temporal model. In the spatial model, the Gaussian process $b(t,s)$ varies over space only and is constant over time. It is defined by a spatial covariance function $\text{Cov}(b(t,s),b(t^{\prime},s^{\prime}))=c_{\theta}(s,s^{\prime})$, $s,s^{\prime}\in\mathcal{D}$, which depends on a set of parameters $\theta\in\Theta\subset\mathbb{R}^{q}$. For more information on covariance functions in general, see williams2006gaussian, and for spatial covariance functions in particular, see cressie2015statistics. In the application below, we use a Matérn covariance function:
[…]
where $\sigma_{1}^{2}$ is a marginal variance parameter, $\rho_{s}$ is a spatial range, or scale, parameter, $\nu$ is a smoothness parameter, $\Gamma(\cdot)$ is the Gamma function, and $K_{\nu}(\cdot)$ is the modified Bessel function of the second kind. For spatio-temporal models, the Gaussian process varies over both space and time, and it is defined by a space-time covariance function $\text{Cov}(b(t,s),b(t^{\prime},s^{\prime}))=c_{\theta}((t,s),(t^{\prime},s^{\prime}))$, $(t,s),(t^{\prime},s^{\prime})\in[0,T]\times\mathcal{D}\subset\mathbb{R}^{3}$. For more information on space-time covariance functions, we refer to cressie2011statistics. In our application, we use an anisotropic spatio-temporal Matérn covariance function:
[…]
where $A=\text{diag}(\rho_{t},\rho_{s},\rho_{s})$ is a diagonal matrix containing a temporal and spatial range parameters $\rho_{t}$ and $\rho_{s}$, respectively. We mainly choose this space-time covariance function since it is a parsimonious model that allows for computationally efficient inference when using a Vecchia approximation and a correlation-based neighbor selection approach (kang2023correlation), see below in Section 3.3 for more details.
@@
$\mathcal{L}(F,\theta)=\int\left(\prod_{i=1}^{N}\prod_{k=0}^{n_{i}-1}\mathcal{L}_{ki}(F(X_{ki}),b(t_{ki},s_{i}))\right)p(b|\theta)db,$ (6)
@@ 2.3 A tree-boosted spatio-temporal Gaussian process model
In the following, we show how to relax the linearity assumption for the fixed effects predictor variable function $F(\cdot)$ using tree-boosting in models with latent Gaussian processes. Tree-boosting (friedman2000additive; friedman2001greedy; buhlmann2007boosting; sigrist2018gradient) is a machine learning technique that often achieves superior prediction accuracy on tabular data sets (nielsen2016tree; shwartz2021tabular; januschowski2022forecasting; grinsztajn2022tree). For instance, recently grinsztajn2022tree showed that tree-boosting outperforms random forest and various state-of-the-art deep neural networks on a large collection of data sets. sigrist2023machine and cheraghali2025sme find similar results for credit risk data.
We assume the model specified in (3) in the subsequent paragraph. In addition, we assume that $F(\cdot)$ is a function in a normed function space $\mathcal{H}$ that is the linear span of a set $\mathcal{S}$ of base learners $f_{j}(\cdot):\mathbb{R}^{p}\rightarrow\mathbb{R}$, which consist of regression trees (breiman1984classification) in this article. Our goal is then to find a joint minimizer for $F(\cdot)\in\mathcal{H}$ and $\theta\in\Theta$ of the functional obtained when plugging $F(\cdot)$ and $\theta$ into the negative log-likelihood:
@@
where $F(X)$ is the row-wise evaluation of $F(\cdot)$ at $X\in\mathbb{R}^{n\times p}$, which is the matrix containing predictor variables for all observations, $n=\sum_{i=1}^{N}n_{i}$, and $\mathcal{L}(F,\theta)$ is given in (6). The minimization of the empirical risk functional in (7) is done iteratively using the latent Gaussian model boosting (LaGaBoost) algorithm given in Algorithm 1, which performs a form of functional gradient descent. In detail, this algorithm iterates between, first, finding a maximum for $\theta$ of $\mathcal{L}(F_{m-1},\theta)$ conditional on the current estimate $F_{m-1}(\cdot)$ and, second, updating the ensemble of trees $F(\cdot)$ using one functional gradient descent step given the current estimates $\theta_{m}$ and $F_{m-1}(\cdot)$. Specifically, the boosting update $f_{m}(\cdot)$ in iteration $m$ is given by the least squares approximation to the vector obtained when evaluating the negative functional gradient of the functional defined in (7) at $(F_{m-1}(\cdot),I_{X_{ki}}(\cdot))$, where $I_{X_{ki}}(\cdot)$ are indicator functions which equal $1$ at $X_{ki}$ and 0 otherwise. Equivalently, $f_{m}(\cdot)$ is the minimizer of a first-order functional Taylor approximation of the functional in (7) with $F(\cdot)=F_{m-1}(\cdot)+f(\cdot)$ around $F_{m-1}(\cdot)$ with an $L^{2}$ penalty on $f(\cdot)$ evaluated at $(X_{ki})$. For more details on the LaGaBoost algorithm, e.g., the calculation of functional gradients; see sigrist2022gaussian; sigrist2023latent.
@@
In order that we can do estimation and prediction in a computationally feasible manner on large data sets in practice, we need to apply some approximations for both the linear Gaussian process and the tree-boosted Gaussian process models. In the following, we describe these approximations. First, the integral in (6) cannot be calculated in closed form, and we approximate it using a Laplace approximation. Laplace approximations are computationally efficient, have asymptotic convergence guarantees, and are accurate for large data sets; see, e.g., kundig2024iterative. Furthermore, to ensure that computations with Gaussian processes scale to large data sets, we use Vecchia approximations (vecchia1988estimation; datta2016hierarchical; katzfuss2017general) for the latent Gaussian process $b$. In spatial statistics, Vecchia approximations have recently “emerged as a leader among the sea of approximations” (guinness2019gaussian) and are often considered as “the most promising class of approximations” (kang2023correlation; rambelli2025accuracy). In brief, Vecchia approximations calculate an approximate sparse reverse Cholesky factor of the precision matrix of the latent Gaussian process $b$ by using an ordered conditional approximation for the density of the Gaussian process. In doing so, every row of this sparse Cholesky factor contains maximally $\tilde{m}$ non-zero entries. For prediction, a Vecchia approximation is applied to the joint distribution of a latent Gaussian process at the training and prediction points. Posterior predictive distributions for the latent Gaussian process are then obtained as conditional distributions of this approximated joint distribution. Further, predictive probabilities for the observable response variables are calculated by numerically integrating over the posterior predictive distribution of the latent Gaussian variable using, e.g., adaptive Gauss-Hermite quadrature (liu1994note). We describe the construction of Vecchia approximations for the latent Gaussian process in more detail in Appendix A.1. For more information on Vecchia-Laplace approximations, including estimation and prediction, see sigrist2023latent and kundig2024iterative. Additionally, for increased computational efficiency, we use the iterative methods of kundig2024iterative for estimation and prediction with Vecchia-Laplace approximations instead of Cholesky decompositions.
@@ 3.2 Predictor variables
From the data provided by Freddie Mac, we construct several static and temporally varying predictor variables. The latter predictor variables are based on information available at the beginning of each year and include the age of the loan in months, the current loan-to-value ratio, and the interest rate spread. The current loan-to-value ratio is calculated by dividing the current unpaid principal balance by the appraised value of the property at the time of origination. To calculate the interest rate spread, we follow calabrese2024impacts and subtract from the loan interest rates the average interest rate currently issued by Freddie Mac for 30-year fixed-rate mortgages. This average interest rate is retrieved from the Federal Reserve Bank of St. Louishttps://fred.stlouisfed.org/series/MORTGAGE30US. Static predictor variables include the borrower’s credit score at the time when the mortgage originated, additional financial indicators, and several characteristics about the property and the mortgage. These predictor variables are commonly chosen when Freddie Mac’s loan-level data set is used in the credit risk literature; see, e.g., hu2019joint and medina2023joint. In addition, we include a number of macroeconomic predictor variables at state and national levels, which vary over time and are primarily sourced from the U.S. Bureau of Economic Analysishttps://bea.gov. A description of all predictor variables is provided in Table 1. Missing values for categorical predictor variables are imputed using the most frequent category, and numeric predictor variables are imputed using the mean. Summary statistics of the numeric and categorical variables are provided in Tables 4 and 5, respectively, in Appendix A.3. For all linear models, we additionally include an intercept term in the predictor variables $X_{ki}$.